# Multimodal Fusion with Relational Learning for Molecular Property Prediction

### v1

@@ 2 Preliminaries
Directed Message Passing Neural Network (DMPNN). The Message Passing Neural Network (MPNN)  (Gilmer et al., 2017) is a GNN model that processes an undirected graph $G$ with node (atom) features $x_{v}$ and edge (chemical bond) features $e_{vw}$. It operates through two distinct phases: a message passing phase, facilitating information transmission across the molecule to construct a neural representation, and a readout phase, utilizing the final representation to make predictions regarding properties of interest. The primary distinction between DMPNN and a generic MPNN lies in the message passing phase. While MPNN uses messages associated with nodes, DMPNN crucially differs by employing messages associated with directed edges  (Yang et al., 2019). This design choice is motivated by the necessity to prevent totters  (Mahé et al., 2004), eliminating messages passed along paths of the form $v_{1}v_{2}\dots v_{n}$, where $v_{i}=v_{i+2}$ for some $i$, thereby eliminating unnecessary loops in the message passing trajectory.
@@
Relational Learning. Original Relation Learning  (Zheng et al., 2021) ensures that different augmented views of the same instance from computer vision tasks share similar features, while allowing for some variability. This approach captures the essential characteristics of the instance, promoting consistency across the views without requiring them to be identical. By doing so, it enhances the model’s ability to generalize and recognize underlying patterns in the data. Suppose $z_{i}$ is the original embdding for the $i-th$ instance. Then $z_{i}^{1}$ is the embedding of first augmented view for $z_{i}$, and $z_{i}^{2}$ is the embedding of second augmented view for $z_{i}$. In this case, the Loss of Relational Learning (RL) is formulated as following:
@@
Multi-Modality Fusion. Multi-Modality Fusion combines diverse heterogeneous data (e.g. text, images, graph) to create a more comprehensive understanding of complex scenarios  (Lahat et al., 2015; Khaleghi et al., 2013; Poria et al., 2015; Ramachandram & Taylor, 2017; Pawłowski et al., 2023). This approach leverages the strengths of each modality, potentially improving performance in tasks like sentiment analysis or medical diagnosis. While challenging to implement due to the need to align different data streams, successful fusion can provide insights beyond what’s possible with single modalities, advancing AI and data-driven decision-making. In particular, the way to fuse different modality should also depends on the dominace of each unimodality  (Pawłowski et al., 2023).
3 Methods
We first explain our proposed modified metric in relational learning to facilitate smooth alignment between graph and referred unimodality. Then, we introduce approaches for integrating multi modalities at different stages of the learning process.
3.1 Modified Relational Learning in pretraining
We propose a modified relational metric by adapting the softmax function as a pairwise weighting mechanism. Let $|\mathcal{S}|$ denote the size of the instance set. The variable ${s}_{i,j}$ represents the learned similarity distribution where $z_{i}$ is the embedding to be trained. On the other hand, ${t}_{i,j}^{R}$ defines the target similarity distribution that captures the relationship between the pair of instances in the given space or modality $R$, where $z_{i}^{R}$ is a fixed embedding. The detailed formulation for the Loss of Modified Relatioal Learning (MRL) is provided below:
@@ 3.1 Modified Relational Learning in pretraining
Notably, unlike other similarity learning approaches  (Wang et al., 2019; Zhang et al., 2021), our method does not rely on the categorization of negative and positive pairs for the pair weighting function. Additionally, the use of the softmax function ensures that the generalized target similarity $t_{i,j}$ adheres to the principles of convergence as following:
@@ 3.2.3 Late Stage: decision-level
4 Experiments
In this section, we begin by presenting the datasets and selected modalities. Subsequently, we showcase the results obtained from MMFRL. Finally, we demonstrate the explainability of the learned molecular representations. (Please refer to the experimental details of pre-training and fine-tuning in the Appendix Section D.)
4.1 Dataset
4.1.1 Selected Modalities for Target Similarity Calculation
Fingerprint: Fingerprints are binary vectors that represent molecular structures, capturing the presence or absence of particular substructures, fragments, or chemical features within a molecule.
NMR (Nuclear Magnetic Resonance): NMR spectroscopy provides detailed insights into the chemical environment of atoms within a molecule. By analyzing the interactions of atomic nuclei with an applied magnetic field, NMR can reveal information about the structure, dynamics, and interactions of molecules, including the connectivity of atoms, functional groups, and conformational changes. In our experiments, $\text{NMR}_{\text{spectrum}}$ provides the information about the overal information of molecule while $\text{NMR}_{\text{peak}}$ provides the information about the individual atoms in the molecule.
4.1.2 Pre-training
4.1.3 Downstream tasks
For Downstream tasks, our model was trained on 11 drug discovery-related benchmarks sourced from MoleculeNet  (Wu et al., 2018a). Eight of these benchmarks were designated for classification downstream tasks, including BBBP, BACE, SIDER, CLINTOX, HIV, MUV, TOX21, and ToxCast, while three were allocated for regression tasks, namely ESOL, Freesolv, and Lipo. The datasets were divided into train/validation/test sets using a ratio of 80%:10%:10%, accomplished through the scaffold splitter  (Halgren, 1996; Landrum, 2006) from Chemprop  (Yang et al., 2019; Heid et al., 2023), like previous works. The scaffold splitter categorizes molecular data based on substructures, ensuring diverse structures in each set. Molecules are partitioned into bins, with those exceeding half of the test set size assigned to training, promoting scaffold diversity in validation and test sets. Remaining bins are randomly allocated until reaching the desired set sizes, creating multiple scaffold splits for comprehensive evaluation.
4.2 Results
4.2.1 The effectiveness of pre-training
We first illustrate the impact of pre-training initialization on performance. As shown in Table 1, the average performance of pre-trained models outperform the non-pre-trained model in all tasks except for Clintox. The results of various downstream tasks indicate that different tasks may prefer different modalities. Notably, the model pre-trained with the NMR modality achieves the highest performance across three classification tasks. Similarly, the model pre-trained with the Image modality excels in three tasks, two of which are regression tasks related to solubility, aligning with findings from prior literature (Xu et al., 2023a). Additionally, the model pre-trained with The fingerprint method achieves the best performance in two tasks, including MUV, which has the largest dataset.
4.2.2 Overall performance of MMFRL
As shown in Table  2 and Table  3, MMFRL demonstrates superior performance compared to all baseline models and the average performance of DMPNN pretrained with extra modalities across all 11 tasks evaluated in MoleculeNet. This robust performance highlights the effectiveness of our approach in leveraging multimodal data. In particular, while individual models pre-trained on other modalities for ClinTox fail to outperform the No-pretraining model (DMPNN), the fusion of these pre-trained models leads to improved performance. Besides, apart from Tox21 and Sider, the fusion models significantly enhances overall performance. In particular, the intermediate fusion model stands out by achieving the highest scores in seven distinct tasks, showcasing its ability to effectively combine features at a mid-level abstraction. the late fusion model achieves the top performance in two tasks. These results underscore the advantages of utilizing various fusion strategies in multimodal learning, further validating the efficacy of the MMFRL framework.
Data Set BBBP bace Sider Clintox HIV MUV Tox21 ToxCast
Data Set ESOL FreeSolv Lipo
AttentiveFP 0.877$\pm$0.029 2.073$\pm$0.183 0.721$\pm$0.001
DMPNN 1.050$\pm$0.008 2.082$\pm$0.082 0.683$\pm$0.016
[…]
GEM 0.798$\pm$0.029 1.877$\pm$0.094 0.660$\pm$0.008
Uni-Mol 0.788$\pm$0.029 1.620$\pm$0.035 0.660$\pm$0.008
MolCLR 1.113$\pm$0.023 2.301$\pm$0.247 0.789$\pm$0.009
$\text{Unimodality}_{avg}$ 0.924$\pm$0.083 1.707$\pm$0.126 0.587$\pm$0.021
$\text{MMFRL}_{early}$ 1.037$\pm$0.170 2.093$\pm$0.090 0.607$\pm$0.034
$\text{MMFRL}_{late}$ 0.763$\pm$0.035 1.741$\pm$0.191 0.525$\pm$0.018
4.3 Analysis of the fusion effect
4.3.1 General Comparison among various Ways of Fusions
In contrast, Late Fusion enables each modality to be explored independently, maximizing the potential of individual modalities without interference from others. This separation allows for a thorough examination of each modality’s contribution. When certain modalities dominate the performance metrics, Late Fusion can maximize on these strengths, ensuring that the most impactful information is utilized effectively. This approach is especially useful in scenarios where the dominance of specific modalities can be leveraged to enhance overall model performance.
4.3.2 Explainability of Learnt Representations
To demonstrate the interpretability of learnt representations of fusion, we present post-hoc analysis for two tasks, ESOL and Lipo, as demonstration. The results showcase learnt representations can capture task-specific patterns and offer valuable insights for molecular design.
ESOL with Intermediate Fusion. As presented in Table 3, the intermediate fusion method  3.2.2 exhibits superior performance on the ESOL regression task for predicting solubility. To further analyze this performance, we employed t-SNE to reduce the dimensionality of the molecule embeddings from 300 to 2, resulting in a heatmap visualized in Figure 2. The embeddings derived from individual modalities prior to fusion do not display a clear pattern, showing no smooth transition from low to high solubility. In contrast, the embeddings by intermediate fusion reveal a distinct and smooth transition in solubility values: molecules with similar solubility cluster together, forming a gradient that extends from the bottom left (indicating lower solubility) to the upper center (representing higher solubility). This trend underscores the effectiveness of the intermediate fusion approach in accurately capturing the quantitative structure-activity relationships for aqueous solubility.
Lipo with Late Fusion. As detailed in Table  3, the Late Fusion method (described in Section  3.2.3) demonstrates superior performance on the Lipo regression task for predicting solubility in fats, oils, lipids, and non-polar solvents. According to Equation  11, the final prediction is determined by the respective coefficients ($w_{i}$) and predictions ($p_{i}$) from each modality.
In Figure 4, we present the distribution of values for the coefficients, predictions, and their products for each modality. Notably, the SMILES and Image modalities display a wide range of values, highlighting their potential to significantly influence the final predictions. This observation aligns with the strong performance achieved when pretraining using either of these two modalities, as shown in Table  1. In contrast, the $\text{NMR}_{\text{Peak}}$ values display a narrower range, indicating its role as a modifier for finer adjustments in the predictions. Furthermore, we observe that the contributions from $\text{NMR}_{\text{Spectrum}}$ and Fingerprint modalities are minimal, with their corresponding values approaching zero. This outcome highlights the advantages of the Late Fusion approach in effectively identifying and leveraging dominant modalities, thereby optimizing the overall predictive performance.
Figure 2: T-SNE visualization depicting the ESOL molecule embeddings for intermediate fusion in Section  3.2.2 alongside molecules within the highlighted region. Each point in the heatmap corresponds to the embeddings of respective molecules in ESOL, with color indicating solubility levels. Red denotes higher solubility, while blue indicates lower solubility. The embeddings derived from individual modalities prior to fusion do not display a clear pattern, the embeddings by intermediate fusion forms a gradient that extends from the bottom left (indicating lower solubility) to the upper center (representing higher solubility).
5 Related work
Contrastive Learning on Molecular Graphs. The primary focus within the domain of contrastive learning applied to molecular graphs centers on 2D-2D graphs comparisons. Noteworthy representative examples: InfoGraph  (Sun et al., 2019) maximizes the mutual information between the representations of the graph and its substructures to guide the molecular representation learning; GraphCL  (You et al., 2020), MoCL  (Sun et al., 2021), and MolCLR  (Wang et al., 2022b) employs graph augmentation techniques to construct positive pairs; MoLR  (Wang et al., 2022a) establishes positive pairs with reactant-product relationships. In addition to 2D-2D graph contrastive learning, there are also noteworthy efforts exploring 2D-3D and 3D-3D contrastive learning in the field. 3DGCL  (Moon et al., 2023) is 3D-3D contrastive learning model, establishing positive pairs with conformers from the same molecules. GraphMVP  (Liu et al., 2022b), GeomGCL  (Li et al., 2022), and 3D Informax  (Stärk et al., 2022) proposes 2D–3D view contrastive learning approaches. To conclude, 2D-2D and 3D-3D comparisons are intra-modality contratsive leraning, as only one graph encoder is employed in these studies. And these approaches often focus on the motif and graph levels, leaving atom-level contrastive learning less explored.
Similarity Learning. Instance-wise discrimination, a crucial facet of similarity learning, involves evaluating the similarity between instances directly based on their latent representations or features  (Wu et al., 2018b). Naive instance-wise discrimination relies on pairwise similarity, leading to the development of contrastive loss  (Hadsell et al., 2006). Although there are improved loss functions such as triplet loss  (Hoffer & Ailon, 2015), quadruplet loss  (Law et al., 2013), lifted structure loss  (Oh Song et al., 2016), N-pairs loss  (Sohn, 2016), and angular loss  (Wang et al., 2017), these methods still fall short in thoroughly capturing relationships among multiple instances simultaneously  (Wang et al., 2019). To address this limitation, a joint multi-similarity loss has been proposed, incorporating pair weighting for each pair to enhance instance-wise discrimination  (Wang et al., 2019; Zhang et al., 2021). Notably, it is crucial to emphasize that employing these pair weightings requires the manual categorization of negative and positive pairs, as distinct weights are assigned to losses based on their categories.
6 Discussion
Accessibility
The code and dataset will be made available upon the date of publication.
@@ C.1 Encoders & Packages
To derive the target similarities, we need to reply on pre-trained encoders or well-defined packages as follows:
@@ C.2 Target Similarity at Graph Level
where $ppm_{l}$ and $ppm_{m}$ are the positions of NMR peaks for the $l^{th}$, $m^{th}$ Carbon atom, $\tau_{1}$ and $\tau_{2}$ are temperature hyper-parameter.
@@ C.4 Configuration of Early Fusion
where $t^{SM}_{i,j}$ denotes the similarity based on SMILES, $t^{C}_{i,j}$ denotes the similarity with respect to 13C NMR spectrum, $t^{I}_{i,j}$ denotes the similarity regarding images, $F$ denotes the similarity based on fingerprints, and $P$ denotes the similarity based on fingerprints. $w_{SM}$, $w_{C}$, $w_{I}$, and $w_{F}$ are the pre-defined weights for their respective similarity, and $w_{SM}+w_{{C}}+w_{{I}}+w_{{F}}+w_{{P}}=1$.
@@ D.1 Pre-Training Setting
During pretraining, we utilized an Adam optimizer with a learning rate set to 0.001, spanning 200 epochs and employing a batch size of 256. The model was trained on around 25,000 data points. The NMR data were experimental data, extracted from NMRShiftDB2  (Steinbeck et al., 2003). Other chemical modalities, such as images, fingerprints and graphs, were produced from SMILES by RDKit  (Landrum, 2006).
@@ D.2.3 Evaluation
To assess the effectiveness of our fine-tuned model, we measure the ROC-AUC for classification downstream tasks, and the root mean squared error (RMSE) metric for regression tasks. In order to ensure a fair and robust comparisons, we conduct three independent runs using three different random seeds for scaffold splitting across all datasets. The reported performance metrics are then averaged across these runs, and the standard deviation is computed as prior works.

### v2

@@ 5.1 Molecular representation with DMPNN
The Message Passing Neural Network (MPNN)  (Gilmer et al., 2017) is a GNN model that processes an undirected graph $G$ with node (atom) features $x_{v}$ and edge (chemical bond) features $e_{vw}$. It operates through two distinct phases: a message passing phase, facilitating information transmission across the molecule to construct a neural representation, and a readout phase, utilizing the final representation to make predictions regarding properties of interest. The primary distinction between DMPNN and a generic MPNN lies in the message passing phase. While MPNN uses messages associated with nodes, DMPNN crucially differs by employing messages associated with directed edges  (Yang et al., 2019). This design choice is motivated by the necessity to prevent totters  (Mahé et al., 2004), eliminating messages passed along paths of the form $v_{1}v_{2}\dots v_{n}$, where $v_{i}=v_{i+2}$ for some $i$, thereby eliminating unnecessary loops in the message passing trajectory.
5.2 Modified Relational Learning in pretraining
@@
Original Relation Learning  (Zheng et al., 2021) ensures that different augmented views of the same instance from computer vision tasks share similar features, while allowing for some variability. Suppose $z_{i}$ is the original embdding for the $i-th$ instance. Then $z_{i}^{1}$ is the embedding of first augmented view for $z_{i}$, and $z_{i}^{2}$ is the embedding of second augmented view for $z_{i}$. In this case, the Loss of Relational Learning (RL) is formulated as following:
@@
We propose a modified relational metric by adapting the softmax function as a pairwise weighting mechanism. Let $|\mathcal{S}|$ denote the size of the instance set. The variable ${s}_{i,j}$ represents the learned similarity where $z_{i}$ is the embedding to be trained. On the other hand, ${t}_{i,j}^{R}$ defines the target similarity that captures the relationship between the pair of instances in the given space or modality $R$, where $z_{i}^{R}$ is a fixed embedding. The detailed formulation for the Loss of Modified Relatioal Learning (MRL) is provided below:
@@
Notably, unlike other similarity learning approaches  (Wang et al., 2019; Zhang et al., 2021), our method does not rely on the categorization of negative and positive pairs for the pair weighting function. Additionally, our use of the softmax function ensures that the generalized target similarity $t_{i,j}$ adheres to the principles of convergence, which results in better ranking consistency between the graph modality and the auxiliary modality, compared with the original Relational Study, as follows:
@@ 5.3.3 Late Stage: decision-level
Data Availability
The pretraining data can be downloaded from NMRShiftDB2. The MoleculeNet dataset is available at MoleculeNet. The DuD-E dataset can be accessed at DuD-E, and the Lit-PCBA dataset can be downloaded from Lit-PCBA.
Code Availability
The code is available in Github: https://github.com/zhengyjo/MMRFL
@@ C.1 Encoders & Packages
To derive the target similarities, we need to reply on pre-trained encoders or well-defined packages as follows in Table C.1:
@@
We selected GIN  Xu et al. (2018) as the graph encoder for Smile, Image, Fingerprint, NMR, and NMR-Peak, respectively. All modalities share a consistent structure, each with 5 layers, an embedding dimension of 128, and a projection dimension of 512. In addition, during pretraining, contrastive learning is performed on each modality independently. For instance, if molecule A only possesses SMILE, IMAGE, and Fingerprint data, but lacks NMR information, it will be included in the training for contrastive learning on SMILE, IMAGE, and Fingerprint, but not on NMR. In contrast, for early fusion, all modalities must be present for the included molecules.
@@ C.2 Target Similarity at Graph Level
where $ppm_{l}$ and $ppm_{m}$ are the positions of NMR peaks for the $l^{th}$, $m^{th}$ Carbon atom, $\tau_{1}$ and $\tau_{2}$ are temperature hyper-parameter. Additionally, we conducted an ablation study to examine the impact of different temperature combinations ($\tau_{1}$ and $\tau_{2}$) on peak self-similarity and its effect on model performance for Atom Alignment with Peak Accuracy. Our findings indicate that the following combinations do not significantly affect performance, as shown in Table C.2. For this analysis, we fixed the GIN depth at 5, set the GIN embedding dimensionality to 128, and kept the projection dimension at 512. The results suggest that the best performance occurs when $\tau_{1}=10^{-5}$ and $\tau_{2}=10^{1}$.
Table C.2: Ablation study of Effect of different $\tau_{1}$ and $\tau_{2}$ combinations on accuracy.
$\tau_{1}$ $\tau_{2}$ Accuracy (%)
$10^{-1}$ $10^{1}$ 89.6
$10^{-1}$ $10^{2}$ 89.8
$10^{-1}$ $10^{3}$ 89.6
$10^{-1}$ $10^{4}$ 88.9
$10^{-1}$ $10^{5}$ 89.3
$10^{-2}$ $10^{1}$ 89.8
$10^{-2}$ $10^{2}$ 89.8
$10^{-2}$ $10^{3}$ 88.8
$10^{-2}$ $10^{4}$ 87.2
$10^{-2}$ $10^{5}$ 89.4
$10^{-3}$ $10^{1}$ 89.2
$10^{-3}$ $10^{2}$ 89.1
$10^{-3}$ $10^{3}$ 89.0
$10^{-3}$ $10^{4}$ 89.7
$10^{-3}$ $10^{5}$ 89.4
$10^{-4}$ $10^{1}$ 89.8
$10^{-4}$ $10^{2}$ 89.7
$10^{-4}$ $10^{3}$ 89.8
$10^{-4}$ $10^{4}$ 89.5
$10^{-4}$ $10^{5}$ 88.4
$10^{-5}$ $10^{1}$ 90.0
$10^{-5}$ $10^{2}$ 89.5
$10^{-5}$ $10^{3}$ 89.6
$10^{-5}$ $10^{4}$ 89.7
$10^{-5}$ $10^{5}$ 89.7
@@ C.4 Configuration of Early Fusion
where $t^{SM}_{i,j}$ denotes the similarity based on SMILES, $t^{C}_{i,j}$ denotes the similarity with respect to 13C NMR spectrum, $t^{I}_{i,j}$ denotes the similarity regarding images, $F$ denotes the similarity based on fingerprints, and $P$ denotes the similarity based on fingerprints. $w_{SM}$, $w_{C}$, $w_{I}$, and $w_{F}$ are the pre-defined weights for their respective similarity, and $w_{SM}+w_{{C}}+w_{{I}}+w_{{F}}+w_{{P}}=1$. The ablation study about how the weight combinations influence the performance of early fusion is shown in Table C.3
Table C.3: Performance of different modality weight combinations across datasets in early fusion.
Modality Weight Dataset
Smiles NMR Image FP Peak BBBP BACE SIDER CLINTOX HIV MUV TOX21 TOXCAST ESOL FREESOLV LIPO
1 0 0 0 0 92.9$\pm$1.5 90.9$\pm$3.3 64.9$\pm$0.3 78.2$\pm$1.9 83.3$\pm$1.1 80.1$\pm$2.5 85.7$\pm$1.2 70.5$\pm$2.5 0.811$\pm$ 0.109 1.623$\pm$ 0.168 0.539$\pm$ 0.017
0 1 0 0 0 91.0$\pm$2.0 93.2$\pm$2.7 68.1$\pm$1.5 87.7$\pm$6.5 80.9$\pm$5.0 80.9$\pm$5.0 85.1$\pm$0.4 71.1$\pm$0.8 0.844$\pm$ 0.123 2.417$\pm$ 0.495 0.609$\pm$ 0.031
0 0 1 0 0 93.1$\pm$2.4 92.9$\pm$1.8 65.3$\pm$1.5 86.2$\pm$6.5 82.3$\pm$0.6 78.7$\pm$1.7 86.0$\pm$1.0 71.0$\pm$1.6 0.761$\pm$ 0.068 1.648$\pm$ 0.045 0.537$\pm$ 0.005
0 0 0 1 0 92.9$\pm$2.3 91.7$\pm$3.6 65.6$\pm$0.7 87.5$\pm$6.0 81.2$\pm$2.5 82.9$\pm$3.1 85.3$\pm$1.3 70.0$\pm$1.4 0.808$\pm$ 0.071 1.437$\pm$ 0.134 0.565$\pm$ 0.017
0 0 0 0 1 93.4$\pm$2.7 89.3$\pm$1.7 62.8$\pm$2.1 86.1$\pm$5.4 82.1$\pm$0.4 75.4$\pm$5.2 84.9$\pm$1.0 70.6$\pm$0.8 0.924$\pm$0.083 1.707$\pm$0.126 0.587$\pm$0.021
0.2 0.2 0.2 0.2 0.2 91.6$\pm$5.0 94.3$\pm$2.4 66.4$\pm$1.9 85.3$\pm$6.8 82.0$\pm$4.2 80.6$\pm$3.2 85.2$\pm$0.6 69.8$\pm$1.1 1.037$\pm$0.090 2.093$\pm$0.090 0.607$\pm$0.034
@@ D.1 Pre-Training Setting
During pretraining, we utilized an Adam optimizer with a learning rate set to 0.001, spanning 200 epochs and employing a batch size of 256. The model was trained on around 25,000 data points. The NMR data were experimental data, extracted from NMRShiftDB2  (Steinbeck et al., 2003). Other chemical modalities, such as images, fingerprints and graphs, were produced from SMILES by RDKit  Landrum (2006).
@@ D.2.3 Evaluation
To assess the effectiveness of our fine-tuned model, we measure the ROC-AUC for classification downstream tasks, and the root mean squared error (RMSE) metric for regression tasks. In order to ensure a fair and robust comparisons, we conduct three independent runs using three different random seeds for scaffold splitting across all datasets. The reported performance metrics are then averaged across these runs, and the standard deviation is computed as prior works. In paritcular, the random selected seeds for respective experiments are drawn from the range between 0 and 20.